How do I find the six trig functions of 9pi/2
Answer by lwsshak3(11628)

How do I find the six trig functions of 9pi/2
***
9π/2=rotating 2 revolutions counter-clockwise (4π)+π/2 which leaves you terminated at π/2
..
sin π/2=1
cos π/2=0
tan π/2=sin/cos=1/0=undefined
csc π/2=1/sin=1/1=1
sec π/2=1/cos=1/0=undefined
cot π/2=cos/sin=0/1=0